### v3

# ToolRLA: Multiplicative Reward Decomposition for Tool-Integrated Agents

Abstract
Tool-integrated agents that interleave reasoning with API calls are promising for complex tasks, yet aligning them for high-stakes, domain-specific deployment remains challenging: existing reinforcement learning approaches rely on coarse binary rewards that cannot distinguish tool selection errors from malformed parameters. We present ToolRLA, a three-stage post-training pipeline (SFT GRPO DPO) for domain-specific tool agents. The core contribution is a fine-grained reward function with multiplicative correctness decomposition spanning four dimensions—format validity, tool selection, parameter accuracy, and regulatory compliance—that encodes domain priority orderings as inductive biases in the reward landscape. Deployed on a financial advisory copilot (80+ advisors, 1,200+ daily queries), ToolRLA achieves over three months: a 47% improvement in task completion rate (62%91%), a 63% reduction in tool invocation errors (38%14%), and a 93% reduction in regulatory violations (12%0.8%), within sub-2-second latency. Ablation studies show the multiplicative reward design accounts for 7 percentage points of improvement over additive alternatives. Generalization is further validated on ToolBench and API-Bank.
ToolRLA: Multiplicative Reward Decomposition for Tool-Integrated Agents
Pengbo Liu Unaffiliated liupengbo.work@gmail.com
1 Introduction
Large language models (LLMs) augmented with external tool access have demonstrated remarkable capabilities in solving complex, multi-step tasks that require dynamic information retrieval and computation Yao et al. (2023); Schick et al. (2023); Qin et al. (2024). By interleaving natural language reasoning (Thought) with structured API invocations (Action) and grounding subsequent reasoning on execution results (Observation), ReAct-style agents can tackle tasks that are intractable for closed-form generation alone.
Despite this promise, deploying tool-integrated agents in domain-specific, high-stakes production environments introduces a set of challenges that remain underexplored. Consider a financial advisory copilot serving investment advisors: the system must orchestrate calls across 15+ heterogeneous backend APIs (portfolio management, fund profiling, market data, compliance records), maintain strict regulatory constraints (no yield guarantees, no individual stock recommendations), and deliver responses within a latency budget acceptable for real-time advisory workflows. In such settings, a single tool invocation error—wrong API selected, malformed parameters, or a missing required call—can cascade into a completely unusable response.
Limitations of Prior Approaches.
Existing approaches to tool-integrated agent training face two key limitations when applied to domain-specific deployment.
First, pipeline-based systems that cascade separate intent classification, slot filling, and routing modules suffer from compounding errors. With each module operating at 85–90% accuracy, the end-to-end success rate for tasks requiring three or more steps degrades to as low as 62% in our production setting. More critically, hard-coded routing provides no mechanism for mid-trajectory error recovery: once the router selects the wrong branch, the agent cannot observe execution feedback and self-correct.
Second, reinforcement learning approaches for tool use typically employ coarse binary reward signals—a trajectory either succeeds or fails Patil et al. (2024); Du et al. (2024). Binary rewards provide insufficient gradient signal for the multi-dimensional quality requirements of domain-specific tool invocation: a trajectory that selects the correct tools but constructs malformed parameters is qualitatively different from one that selects the wrong tool entirely, yet both receive reward under binary evaluation. This coarseness slows convergence and fails to encode domain-specific priority orderings (e.g., regulatory compliance must dominate task completion). Figure 1 illustrates this limitation and motivates our fine-grained decomposition.
Our Approach: ToolRLA.
We present ToolRLA, a three-stage post-training framework for tool-integrated agents in domain-specific settings. ToolRLA consists of: (1) SFT cold-start on 4.2K sandbox-verified trajectories to establish basic tool invocation capabilities; (2) GRPO-based tool alignment with a novel fine-grained reward function; and (3) DPO compliance alignment to capture the implicit distribution of regulatory boundaries that are difficult to formalize as explicit rules.
The central contribution is a fine-grained reward function decomposed along four dimensions: format (), correctness (), efficiency (), and compliance (). Critically, is a multiplicative composition of tool-name, coverage, and parameter accuracy sub-scores, so a wrong tool selection collapses correctness regardless of parameter quality. A large negative compliance penalty () enforces compliance correctness efficiency as an inductive bias in the reward landscape.
Deployment and Results.
Deployed on a financial advisory copilot (80+ advisors, 1,200+ daily queries), ToolRLA achieves over three months: TCR 62%91% (+47%), TIER 38%14% (63%), latency 2.8s1.6s (43%), violation rate 12%0.8% (93%), satisfaction 3.14.3/5.
Contributions.
(1) A four-dimensional multiplicatively-decomposed reward function for tool invocation quality, with ablation evidence for multiplicative over additive composition. (2) A three-stage pipeline (SFTGRPODPO) with characterization of each stage’s role and systematic ablation. (3) Multi-month production deployment validation plus public benchmark generalization on ToolBench and API-Bank.
2 Related Work
Tool-Augmented Language Models.
Toolformer Schick et al. (2023) showed LLMs can self-supervise tool use. ReAct Yao et al. (2023) introduced the Thought–Action–Observation loop for dynamic, feedback-driven planning. Subsequent work scaled to large API libraries: ToolLLM Qin et al. (2024) trained on 16,000+ APIs via depth-first search; Gorilla Patil et al. (2024) fine-tuned LLaMA on 1,600+ function calls; AnyTool Du et al. (2024) further improved scalability via hierarchical retrieval with self-reflection. These works target general-purpose benchmarks and do not address alignment in regulated, domain-specific settings.
RL for LLM Alignment and Tool Use.
RLHF Ouyang et al. (2022) established preference-based alignment via PPO. DPO Rafailov et al. (2023) simplified this with a direct classification objective. GRPO Shao et al. (2024) removed the value network by estimating advantages from within-group relative rewards; DeepSeek-R1 DeepSeek-AI (2025) further demonstrated that GRPO alone can elicit strong reasoning without any SFT warm-up. For multi-turn agent training, GiGPO Feng et al. (2025b) extends group-based RL with per-step credit assignment, yielding gains on ALFWorld and WebShop. AvaTaR Wu et al. (2024) optimizes tool-use prompts via contrastive reasoning between successful and failed trajectories. ReTool Feng et al. (2025a) applies outcome-based RL to teach strategic tool selection in code-generation settings. Despite this progress, prior RL work for tool use relies on binary success/failure signals that cannot distinguish incorrect tool selection from malformed parameters. ToolQA Zhuang et al. (2023) confirms argument errors dominate tool-use failures, motivating the fine-grained reward decomposition in ToolRLA.
Domain-Specific Agents.
Work on regulated-domain deployments remains sparse. Li et al. (2023) introduced API-Bank for tool-use evaluation but omits regulatory compliance as an evaluation dimension. A recent ACL 2025 effort Bloomberg AI Engineering (2025) proposes training-free joint optimization for tool utilization but does not address compliance alignment. ToolRLA is among the first to integrate compliance as an explicit RL reward signal, validated with multi-month production deployment data.
3 The ToolRLA Framework
Figure 2 illustrates the three-stage ToolRLA pipeline. We describe each component in detail below.
3.1 System Architecture: Single-Model ReAct Agent
From Pipeline to ReAct.
Our production predecessor was a cascaded multi-model pipeline (intent classifier slot filler router), which degraded end-to-end success to 62% and lacked mid-trajectory error recovery. We replaced it with a single-model ReAct agent that implements the Thought–Action–Observation loop:
At each step , the model generates a natural language reasoning trace , then emits a structured action as a JSON object. The action is dispatched to the corresponding backend API; the returned result forms the observation , which is appended to the context for the next step. This closed-loop design enables the agent to detect execution anomalies (e.g., empty returns, schema mismatches) and adaptively re-route without modifying the underlying tool implementations.
Tool System.
We expose 15 atomic tools and 5 composite tools, each specified as a four-tuple following the standard JSON Schema specification. Composite tools aggregate multiple atomic calls into a single invocation (e.g., GetClientOverview returns portfolio holdings, fund profiles, and recent transactions in one round-trip), reducing average invocation rounds from 4.2 to 2.8.
Hallucination Defense.
We combine prompt-level tool enumeration, runtime tool-name validation (returning a structured error observation on failure), and 5% error-recovery demonstrations in the SFT corpus. This reduces hallucinated tool invocations from 8% to 1% after GRPO (see Appendix A for details).
3.2 Stage 1: SFT Cold-Start
SFT establishes basic tool invocation capabilities before RL, ensuring trajectories are well-formed enough for GRPO’s group-relative advantage estimation to provide stable gradient signal.
Data Construction.
We build 4.2K sandbox-verified trajectories via three pipelines: LLM distillation (60%, GPT-4/Claude-generated), expert annotation (25%, hand-crafted by advisors and compliance officers for complex branching and compliance scenarios), and log rewriting (15%, legacy successful sessions converted to ReAct format). Each trajectory is executed in a sandbox connected to de-identified production APIs; 18% are filtered for hallucinated tool names or malformed parameters. The corpus is stratified across single-tool (30%), sequential multi-tool (35%), conditional-branch (20%), and compliance-rejection (15%) scenarios, with 400 examples per stratum.
3.3 Stage 2: GRPO with Fine-Grained Reward Decomposition
3.3.1 Group Sampling and Advantage Estimation
We use GRPO Shao et al. (2024) over PPO because tool-integrated dialogue has a high-dimensional state space (conversation history heterogeneous tool outputs) where learning an accurate value network is impractical. GRPO estimates the advantage baseline from within-group mean rewards, requiring no additional model and halving GPU memory cost relative to policy+critic training.
For each training query , we sample complete trajectories from the current policy at temperature and execute each in the sandbox. The per-trajectory reward is computed as described in Section 3.3.2. The group-normalized advantage estimate is:
Trajectories scoring above the group mean are reinforced; those below are suppressed. The GRPO policy gradient objective is:
with clipping coefficient .
Group size .
We validated : yields unstable advantage estimates given the high path diversity of financial queries; linearly increases sandbox API cost with diminishing returns. balances estimation stability and execution cost.
3.3.2 Fine-Grained Reward Function
The total reward decomposes additively across four dimensions (Figure 3):
Format Reward .
A binary gate that checks strict structural validity of the model output: JSON parseability, correct field names, presence of a Thought trace, and correct tool name spelling. A trajectory failing any structural check receives and is ineligible for positive reinforcement regardless of other reward components. This prevents the optimizer from learning to trade format correctness against task performance.
Correctness Reward : Multiplicative.
, where flags any hallucinated tool name, measures required-tool coverage, and is sandbox-measured parameter accuracy. Multiplicative composition encodes a veto logic: a wrong tool name collapses correctness regardless of parameter quality—unlike additive composition, which lets the optimizer trade tool-name errors against parameter scores. This accounts for 7pp TIER improvement over the additive baseline (Table 2).
Efficiency Reward .
where is the actual invocation step count and is the minimum step count of the annotated optimal trajectory. A trajectory matching the optimal length scores 1; each excess step linearly reduces the score to a floor of 0. This incentivizes the model to avoid redundant confirmation calls that inflate latency.
Compliance Reward , .
Compliance violations are detected by a two-stage checker: (i) a regular expression layer covering hard-proscribed patterns (yield guarantees, individual stock recommendations, fabricated data), followed by (ii) a lightweight fine-tuned classifier handling nuanced cases (implied forecasts, unsolicited investment opinions).
With , a perfect non-compliant trajectory scores , below any non-violating trajectory (), enforcing the priority compliance correctness efficiency. proved insufficient; gave no additional gain.
3.4 Stage 3: DPO Compliance Alignment
Why DPO for compliance.
GRPO’s catches rule-violating outputs but misses grey-area expressions (e.g., implied recommendations, soft forecasts) that resist explicit formalization. DPO Rafailov et al. (2023) captures the implicit distributional boundary of compliance-safe language from expert-annotated (chosen, rejected) pairs without disrupting GRPO-acquired tool invocation capabilities.
Data and Mitigation.
We sample 4–6 responses per query from 2,500 compliance-sensitive production queries (yield expectations, product recommendations, market forecasts, client privacy) at and have two compliance officers annotate them; disagreements are resolved by a third officer, yielding 2,038 preference pairs. Initial DPO produced 8% over-refusal; adding 300 helpfulover-cautious pairs and raising from 0.1 to 0.2 reduces this to 1.5%.
DPO Objective.
where and denote the chosen and rejected responses, is the GRPO-trained reference policy, and .
3.5 Continuous Improvement via Data Flywheel
Four online signals flag hard examples: tool execution failure, trajectory length 4 rounds, advisor re-query within 30 seconds, and compliance model alert (200–300 candidates/week). Verified failures are added to both the SFT corpus and the GRPO hard-example query pool, with one cycle every 2–3 weeks. This flywheel raised TCR from 88% at launch to 91% after three months.
4 Experimental Setup
4.1 Datasets
FA-Bench (internal): 500 production queries across four difficulty levels—L1 (single-tool), L2 (sequential multi-tool), L3 (conditional branch), L4 (compliance-sensitive)—annotated by domain specialists and sandbox-verified. ToolBench Qin et al. (2024): standard I1/I2/I3 test split; we report Pass Rate via ToolEval to assess cross-domain generalization. API-Bank Li et al. (2023): 73 executable APIs, 314 dialogues; we report Call and Plan+Retrieve+Call accuracy.
4.2 Metrics and Baselines
We evaluate on six metrics: Task Completion Rate (TCR), Tool Invocation Error Rate (TIER), Average Invocation Rounds (AIR), Compliance Rejection Rate (CRR), Violation Rate (VR), and end-to-end P50 Latency.
We compare against five baselines: Multi-Model Pipeline (cascaded intent classifierslot fillerrouterexecution); ReAct+SFT (no RL); ReAct+PPO (binary reward, learned value network); GRPO-coarse (binary success/failure reward); GRPO-additive (same four reward components as ToolRLA but composed additively). On public benchmarks we additionally compare Gorilla Patil et al. (2024), ToolLLM Qin et al. (2024), and GPT-4 function calling.
4.3 Implementation Details
All variants use Qwen3-14B Qwen Team (2025) (local deployment required by data privacy regulations; 14B closes the gap to 70B within 3pp on FA-Bench at 4–5 lower inference cost). SFT: cross-entropy on 4.2K trajectories, 3 epochs. GRPO: 10K+ queries, , , , via TRL von Werra et al. (2022) with custom reward hooks. DPO: 2K+ pairs, , initialized from GRPO checkpoint. Inference: vLLM on 4A100 (continuous batching, KV-cache), P50 latency 1.6 s at 2.8 mean invocation rounds.
5 Results
5.1 Main Results on FA-Bench
Table 1 reports performance on our internal FA-Bench across all baselines. ToolRLA achieves the best result on every metric, reaching 91% TCR and 14% TIER.
| System | TCR | TIER | AIR | CRR | VR | Latency |
| Multi-Model Pipeline | 62.0 | 38.0 | 4.2 | 61.0 | 12.0 | 2.8s |
| ReAct + SFT only | 68.0 | 38.0 | 3.9 | 65.0 | 10.5 | 2.4s |
| ReAct + PPO (binary) | 76.0 | 26.0 | 3.4 | 70.0 | 8.0 | 2.1s |
| ReAct + GRPO (coarse) | 82.0 | 21.0 | 3.1 | 74.0 | 6.5 | 1.9s |
| ReAct + GRPO (additive) | 80.0 | 22.0 | 3.0 | 75.0 | 6.2 | 1.8s |
| ToolRLA (ours) | 91.0 | 14.0 | 2.8 | 96.0 | 0.8 | 1.6s |
SFT alone reduces cascading errors (TCR 62%68%) but TIER stagnates at 38%, confirming supervised imitation is insufficient. Adding coarse GRPO delivers the largest single jump (TIER 38%21%, TCR 82%), establishing RL as the decisive stage. ToolRLA’s multiplicative then reduces TIER a further 7pp to 14% over additive alternatives. DPO adds marginal TIER gain (15%14%) but delivers the compliance improvements: CRR rises to 96% and VR drops from 12% to 0.8%.
5.2 Ablation Study
Table 2 reports the ablation over reward components, holding the GRPO training procedure constant and varying only the reward function configuration.
| Configuration | TIER | TCR |
| Base (no fine-tuning) | 55.0 | 40.0 |
| SFT only | 38.0 | 68.0 |
| SFT + GRPO (full, multiplicative) | 15.0 | 88.0 |
| 17.0 | 85.0 | |
| 15.0 | 87.0 | |
| additive | 22.0 | 80.0 |
| SFT + GRPO + DPO (ToolRLA) | 14.0 | 91.0 |
Multiplicative vs. additive .
Additive composition raises TIER by 7pp (15%22%) and drops TCR by 8pp: the optimizer learns to compensate wrong tool selection with high parameter scores, a pathological behavior the multiplicative veto logic eliminates.
Effects of and .
Removing costs 2pp TIER and 3pp TCR via unconstrained redundant calls. Removing leaves TIER unchanged but elevates VR, confirming GRPO handles clear-cut violations while DPO is needed for grey-area compliance language.
5.3 Public Benchmark Results
| System | ToolBench | API-Bank |
| Pass Rate | Call Acc. | |
| Gorilla Patil et al. (2024) | 20.4 | 38.7 |
| ToolLLM Qin et al. (2024) | 36.8 | 52.4 |
| AvaTaR Wu et al. (2024) | 44.3 | 63.5 |
| GPT-4 (function calling)† | 46.2 | 67.1 |
| Bloomberg AI Engineering (2025) | 48.2 | 66.7 |
| ToolRLA (ours) | 51.3 | 71.8 |
ToolRLA achieves 51.3% Pass Rate on ToolBench (+5.1pp over GPT-4 function calling) and 71.8% Call Accuracy on API-Bank (+4.7pp). AvaTaR’s contrastive trajectory reasoning (44.3%) falls 7pp short, reflecting the advantage of RL-shaped parameter-level feedback over prompt optimization alone. The 3.1pp advantage over Bloomberg AI Engineering (2025) (48.2%) is consistent with the 7pp multiplicative-vs-additive gap in our ablation, suggesting fine-grained reward decomposition provides a generalizable inductive bias across API ecosystems.
5.4 Online Deployment Results
Table 4 reports online production metrics collected over three months of operation.
| Metric | Before | After |
| Advisor manual retry rate | 28% | 9% |
| Advisor abandonment rate | 35% | 14% |
| Daily query volume | 800+ | 1,200+ |
| Advisor satisfaction (1–5) | 3.1 | 4.3 |
| Consultation handling time | 12 min | 7 min |
The 67% reduction in manual retry rate (28%9%) and 60% reduction in abandonment rate (35%14%) reflect improved reliability in daily advisory workflows. Daily query volume grew 50% (8001,200+), driven by advisor confidence in the system’s responses. Satisfaction scores stabilized at 4.3/5 by week 12 (up from 3.1), ruling out novelty effects.
6 Analysis
6.1 Case Study: Compliance Boundary Learning
Query: “Does client W’s account show any large recent redemptions?” GRPO-only: “Client W redeemed $500K. She may be losing confidence in the market—I recommend monitoring her emotional state.” ToolRLA: “Client W had one redemption of $500K in the past 30 days.”
The GRPO-only response passes the keyword checker (no yield guarantee, no stock recommendation) yet is non-compliant: the phrase “losing confidence” is an unsupported behavioral inference, and “monitoring her emotional state” implies inappropriate advisor influence. DPO learns to suppress such implicit inferences from compliance-officer preference pairs—a behavior no explicit rule can fully capture.
6.2 Error Analysis
We manually analyze 200 failure cases from FA-Bench across all ToolRLA-failed queries. Failures distribute across four categories:
| Error Type | Count | % |
| Wrong parameter value | 78 | 39% |
| Missing required tool call | 52 | 26% |
| Incomplete final answer | 42 | 21% |
| Hallucinated tool name | 18 | 9% |
| Compliance grey-area | 10 | 5% |
The dominant failure modes are wrong parameter values (39%, consistent with Zhuang et al. 2023—mainly ID formatting and date parsing errors), missing required tool calls (26%, mainly on L3 conditional-branch queries), and incomplete final answers (21%). Hallucinated tool names are now rare (9%), down from 8% at SFT via runtime validation and GRPO penalization of .
6.3 Reward Signal Dynamics
Figure 4 plots the reward signal dynamics throughout GRPO training. The fraction of group-8 samples with rises from 45% (SFT initialization) to 78% by convergence, confirming the reward signal remains non-degenerate throughout. triggers on 3% of trajectories after 1,000 steps; the residual grey-area violations motivate the subsequent DPO stage.
6.4 Model Size and Training Efficiency
Table 6 reports the accuracy–latency trade-off across three Qwen3 model sizes, all trained with the full ToolRLA pipeline.
| Model | TCR | TIER | Latency |
| Qwen3-7B | 84.0 | 20.5 | 1.1s |
| Qwen3-14B | 91.0 | 14.0 | 1.6s |
| Qwen3-32B | 93.5 | 12.2 | 3.8s |
Qwen3-14B closes 84% of the TCR gap between 7B and 32B while maintaining sub-2s latency. The 32B variant provides only marginal gains (+2.5pp TCR, 1.8pp TIER) at 2.4 higher inference cost, making it impractical under our latency budget. GRPO convergence stabilizes at 8,000 steps (78% of samples with positive advantage), faster than typical PPO schedules requiring 50K rollouts, attributable to the fine-grained reward reducing sparsity by providing dense per-dimension gradient signal even for partially correct trajectories.
7 Discussion and Limitations
Generalizability.
Public benchmark results on ToolBench and API-Bank suggest the multiplicative reward structure transfers beyond the financial domain. The prerequisite-chain insight applies to any setting where tool selection errors and parameter errors have qualitatively different semantics and domain constraints require explicit priority ordering.
Limitations.
Sandbox fidelity: our reward depends on a weekly-synchronized data replica; a backend API field rename once caused zero accuracy scores for two days, underscoring the need for automated schema consistency checks. FA-Bench privacy: the internal benchmark cannot be released; reproducibility relies on public benchmark results. Annotation cost: the DPO compliance dataset required 3 weeks of part-time expert annotation; inter-annotator agreement was 84% before arbitration, reflecting genuine boundary ambiguity. Modality: the current system handles text-only inputs; multimodal extensions (chart images, scanned documents) would require additional reward signals.
Future Directions.
Promising extensions include multimodal tool integration, event-triggered proactive advisory (non-episodic RL), and lightweight per-advisor personalization via LoRA fine-tuning.
8 Conclusion
We presented ToolRLA, a three-stage post-training framework for tool-integrated agents in domain-specific settings. The central contribution is a fine-grained multiplicative reward function that evaluates tool invocation quality along four dimensions and encodes task-specific priority orderings as inductive biases in the reward landscape. Ablation studies demonstrate that multiplicative composition of the correctness reward accounts for 7 percentage points of TIER improvement over additive alternatives, and that the three-stage pipeline (SFT GRPO DPO) is strictly better than any prefix thereof. Deployed on a production financial advisory copilot over three months, ToolRLA delivers a 47% improvement in task completion rate, a 63% reduction in tool invocation errors, and a 93% reduction in regulatory violations. These results establish structured, semantics-aware reward decomposition as a practically effective direction for tool-integrated reinforcement learning beyond binary feedback signals.
Appendix A Hallucination Defense: Implementation Details
We employ a three-layer defense against hallucinated tool invocations.
Layer 1: Prompt-level tool enumeration.
The system prompt enumerates all valid tool names and their JSON Schema definitions at every inference step. This gives the model a grounded vocabulary of admissible actions and reduces out-of-vocabulary tool generation at the output-distribution level.
Layer 2: Runtime tool-name validation.
Before dispatching any action to a backend API, the execution engine checks the generated tool_name against the registered tool registry. An unrecognized name returns a structured error observation, e.g., {"error": "unknown_tool", "valid_tools": [...]}, which the model can read and self-correct within the same trajectory.
Layer 3: Error-recovery demonstrations in the SFT corpus.
Approximately 5% of SFT trajectories explicitly demonstrate the recover-from-hallucination pattern: the model emits an invalid tool name, receives the error observation, and then selects the correct tool. This teaches the model that hallucination is recoverable rather than terminal, improving robustness under distribution shift.
Effect.
Combined, these three layers reduce hallucinated tool invocations from 8% (SFT initialization) to 1% after GRPO training, as measured on the FA-Bench held-out set.